A Frisbee is 26.0 cm in diameter and has half its mass concentrated in the rim, the rest spread uniformly over its area. The Frisbee's rotational inertia is 1.67 g⋅m2. What's its mass?
Answer #1

For rotation about the axes perpendicular to the disk and through the center, a uniform disk has moment of inertia 1/2 m R^2 . A ring has moment of inertia m R^2 .

d =26 cm , R=13 cm

Here the m = 1/2 M

I_frisbee = 1/2 m R^2 + m R^2

= 1/2 (1/2 M) R^2 + (1/2 M) R^2

1.67 = ((1/4)M*(0.13*0.13)) +(0.5*M*0.13*0.13)

M =131 g = 0.131 kg